Find less than six times a number is 104. Find the number
Answer by ptaylor(1894)


I AM ASSUMING "FIND" SHOULD BE "FOUR"

Let x=the number
Then 6x=six times the number
And 6x-4=four less than six times the number and we are told this equals 104. So our equation to solve is:
6x-4=104 add 4 to both sides
6x-4+4=104+4 collect like terms
6x=108 divide both sides by 6
x=18-----------the number
ck
6*18-4=104
108-4=104
104=104

Hope this helps----ptaylor